write the equation of a line parallel to y=3/4x+2 and passes through (-4,4)
Answer by sdmmadam@yahoo.com(530)

write the equation of a line parallel to y=3/4x+2 and passes through (-4,4)
The given line is
y=(3/4)x+2 ----(1)
Required to find the equation to the line parallel to (1)
and passing through P(-4,4)
Any line parallel to (1) is given by
y = (3/4)x + k ----(2)
(lines parallel and so slopes equal and hence the equation to a line parallel to y=mx +b is given as y = mx +k)
Given that P(-4,4) is a point on (2)
Therefore putting x=-4 and y=4 in (2)
4=(3/4)X(-4)+k
4=-3+k
4+3 = k
7=k
Therefore k=7
Putting k=7 in (2)
The required line is
y = (3/4)x +7